# envname-P envname#1

### v1

@@ Abstract
The ability of CodeLLMs to generate executable and functionally correct code at the repository-level scale remains largely unexplored. We introduce RepoExec, a novel benchmark for evaluating code generation at the repository-level scale, emphasizing executability and correctness. RepoExec provides an automated system that verifies requirements and incorporates a mechanism for dynamically generating high-coverage test cases to assess the functionality of generated code. Our work explores a controlled scenario where developers specify necessary code dependencies, challenging the model to integrate these accurately. Experiments show that while pretrained LLMs outperform instruction-tuning models in correctness, the latter excel in utilizing provided dependencies and demonstrating debugging capabilities. RepoExec aims to provide a comprehensive evaluation of code functionality and alignment with developer intent, paving the way for more reliable and applicable CodeLLMs in real-world scenarios.
@@ 1 Introduction
Code Large Language Models (CodeLLMs) have emerged as a prominent solution to assist with coding tasks recently wang2021codet5; wang2023codet5+; feng2020codebert; allal2023santacoder; li2023starcoder; lozhkov2024starcoder; guo2024deepseek; pinnaparaju2024stable; zheng2024opencodeinterpreter; roziere2023code; nijkamp2022codegen; luo2023wizardcoder; xu2022systematic; bui2023codetf. While CodeLLMs are adept at capturing the immediate context of the code being written, they often do not fully exploit the broader context available within the entire code repository. This limitation stems from the token length restrictions of general LLMs, which also apply to CodeLLMs. Consequently, suggestions generated by these models may not be optimally aligned with the project’s architecture or intended functionality, as they tend to overlook the rich information embedded in related files, class hierarchies, dependencies, and more. Therefore, evaluating the ability of CodeLLMs on code generation at the repository-level scale is crucial to accurately measure their performance in real-world scenarios. There are existing repository-level code generation benchmarks , such as RepoBench liu2023repobench, RepoCoder zhang2023repocoder, CrossCodeEval DBLP:conf/nips/DingWADTJRNBRX23, CoCoMIC ding-etal-2024-cocomic-code, and DevEval li2024deveval, Although these benchmarks provide a comprehensive perspective on evaluating code generation at the repository-level scale, they lack 3 major features that prevent them from accurately reflecting real-world coding scenarios. First, code generation in the real world requires the code to be executable, and none of the existing benchmarks focus on this aspect. Second, these benchmarks rely on simple methods to retrieve contexts, such as including all imports and function calls (RepoBench) or simply retrieving similar files as contexts (RepoCoder), which might introduce noise into the contexts and impact the performance of code generation systems on these benchmarks. Third, generating real world code at repository-level scale requires the CodeLLMs to be able to capture contexts (both cross-file and in-file effectively), which may pose several questions such as if the dependencies provided by human are really the dependencies that the LLM need, or how well the CodeLLMs process the dependencies given that lots of depdencies feed into the prompt, do they use all of them or skip some of them?
@@ 2 Related works
Coding-related tasks have been crucial for assessing the performance of Large Language Models (LLMs), with code generation emerging as a primary focus (chen2021evaluating; li2023starcoder; jiang2024mixtral; touvron2023llama; roziere2023code; xu2022systematic; allal2023santacoder; nijkamp2022codegen). Early benchmarks have been introduced to address this issue yin2018learning; iyer2018mapping; manh-etal-2023-vault; chen2021evaluating; austin2021program; hendrycksapps2021; however, they often had limited scope or employed weak evaluation approaches. Some benchmarks yin2018learning; iyer2018mapping; manh-etal-2023-vault exhibit domain diversity akin to real-world applications; however, their evaluation methodologies are constrained to match-based metrics, thereby decreasing the reliability of these benchmarks chen2021evaluating. Meanwhile, benchmarks with reliable evaluation approaches such as HumanEval chen2021evaluating, MBPP austin2021program and APPS hendrycksapps2021 often entail limitations to specific domains like competitive programming. Besides, all the benchmarks mentioned primarily focus on standalone function generation, lacking consideration for cross-contextual and dependency invocation scenarios.
Several recent studies have introduced frameworks and benchmarks for repository-level code generation (ding-etal-2024-cocomic-code; DBLP:conf/icml/ShrivastavaLT23; DBLP:conf/nips/DingWADTJRNBRX23; liao2023context; liu2023repobench). These studies closely align with real-world scenarios, underscoring the importance of cross-contextual considerations. However, these works are still limited to match-based evaluation methods. A recent study li2024deveval introduced the DevEval benchmark for code generation within repository contexts, evaluating performance based on extracted tests available in the repository. These benchmarks primarily focus on assessing the functional correctness of generated outputs but have not extensively investigated the correctness in conjunction with the ability to utilize dependency contexts.
@@ 3.1 Data Source
Developing an executable benchmark within repository contexts is challenging due to complex setup requirements and frequent lack of clear installation guidelines in repositories. Previous studies (ding-etal-2024-cocomic-code; DBLP:conf/icml/ShrivastavaLT23; DBLP:conf/nips/DingWADTJRNBRX23; liao2023context) have often relied on match-based metrics for evaluation, which may not fully capture code functionality. In addition, test cases are essential for assessing code functionality. However, extracting test cases from repositories (DBLP:conf/emnlp/ZhangCZKLZMLC23; li2024deveval; zhang2024codeagent) often relies on available functions and heuristic rules, limiting adaptability and excluding data without existing tests. For instance, li2024deveval found that over 99% of functions were discarded due to the absence of suitable tests. We propose a dataset collection pipeline to ensure that repositories can build executable environments and that test cases are automatically generated.
@@ 3.2 Functions and Dependencies Extraction
We extract functions and their dependencies from repositories, considering only those suitable for function-level code generation, akin to benchmarks like HumanEval chen2021evaluating and MBPP austin2021program. Using tree-sitter, we parse files into Abstract Syntax Trees (AST) to extract functions, focusing on those with comprehensive docstrings and excluding functions used as entry points or that do not produce verifiable outputs.
@@ 3.3 Test case generation
To overcome the limitation of requiring available tests in the repository for evaluation (as mentioned in Section 3.1), we leverage large language models (LLMs) to generate test cases automatically. Our proposed approach ensures the correctness of created test cases while also controlling and enhancing test coverage. To execute and validate the generated test cases, it is necessary to configure each repository to create an executable environment. Follow lemieux2023codamosa, we use pipreqs [^1] https://github.com/bndr/pipreqs to identify each project’s dependencies.
@@ 3.3.1 Correctness Control
In this stage, we present the procedure for generating initial test cases and ensure the tests’ correctness. Specifically, we use CodeLlama-13b roziere2023code and provide the model with the prompt detailed in Appendix C. After generation, we extract the first 20 assertion statements to construct the test cases. To ensure their correctness, we employ syntax and execution-based filtering methods to exclude low-quality test cases.
@@ 3.3.2 Coverage Enhancement
Weak unit tests may allow incorrect implementations to pass liu2023is. To address this, we propose a strategy for enhancing test coverage using LLMs. Given the complexity of this task, requiring a strong understanding of code, we utilize GPT-3.5 to improve the quality of test cases. We provide GPT-3.5 with three prompts (see Appendix C) to handle challenging scenarios, including edge and corner cases. The initial generated tests serve as few-shot examples. We then extract the newly generated tests and ensure their correctness using our methodology from Section 3.3.1. Table 1 shows a coverage improvement of about 4% after enhancement, reaching 96.25%. The performance gap has also increased to over 5% (Appendix B), indicating greater robustness in the evaluation.
@@ 4.1 Data Formatting
We follow muennighoff2023octopack to define 2 types of prompt formats in our evaluation of RepoExec across LLMs: (1) BasePrompt, which concatenates all contexts with the target functions (Figure 1), and (2) InstructPrompt, which includes specific instructions for the LLMs to follow, utilizing two variations as input formats (further details and examples in Appendix D).
@@ 4.2 Dataset Stastistic
Dataset #Samples Repo context Test case Coverage (%) CoNaLA yin2018learning 500 - CONCODE iyer2018mapping 2,000 - HumanEval chen2021evaluating 164 ✓, H 99.43 MBPP austin2021program 974 ✓, H 98.48 RepoCoder DBLP:conf/emnlp/ZhangCZKLZMLC23 373 ✓, P - CrossCodeEval DBLP:conf/nips/DingWADTJRNBRX23 2,665 - CoCoMIC ding-etal-2024-cocomic-code 6,888 - DevEval li2024deveval 1,874 ✓, P - RepoExec 355 ✓, A 92.46    + coverage-enhancement 96.25
@@
Table 1 compares the details of RepoExec with those of existing code generation benchmarks. Benchmarks that exclude execution-based evaluation yin2018learning; iyer2018mapping; DBLP:conf/nips/DingWADTJRNBRX23; ding-etal-2024-cocomic-code may gather substantial amounts of data; however, they are inadequate for assessing the quality of the generated code. For HumanEval and MBPP, the majority of problems involve standalone functions, which do not reflect real-world scenarios. Besides, benchmarks that rely on human-annotated and pre-existing test cases chen2021evaluating; austin2021program; zhang2023repocoder; li2024deveval are challenging to scale and control the test coverage. Finally, repository-context benchmarks (RepoCoder zhang2023repocoder, CrossCodeEval DBLP:conf/nips/DingWADTJRNBRX23, CoCoMIC ding-etal-2024-cocomic-code, and DevEval li2024deveval) primarily focus on evaluating retrieval modules. In contrast, our approach emphasizes the generation module’s ability to understand and utilize human-provided dependency contexts.
@@
#No problem #No testcase Avg tokens Cross-file Total Avg Coverage (%) Avg Prompt Solution Full context 22.8% 355 96.25 98.38 362.92 78.46 Medium context - - - - 253.05 - Short context - - - - 179.66 -
@@ 5.1 Evaluation metrics
We use the well-known Pass@k chen2021evaluating to measure the functional correctness of generation outputs. Futhermore, to assess the models’ ability to utilize provided dependencies in accordance with human intent, we introduce the Dependency Invocation Rate (DIR). This metric represents the percentage of invoked dependencies out of the total number of dependencies provided. Let $D_{g}$ denote the set of identifiers in the generated output, and $D_{s}$ denote the set of provided dependencies extracted from the solution. The Dependency Invocation Rate (DIR) is calculated as follows:
@@ 5.2 Evaluation Results
Model Full context Medium context Small context pass@1 pass@5 DIR pass@1 pass@5 DIR pass@1 pass@5 DIR BasePrompt Pre-models CodeLlama-34b-Python roziere2023code 42.93 49.54 68.85 35.92 42.95 58.15 39.80 45.79 64.23 CodeLlama-13b-Python roziere2023code 38.65 43.24 62.26 32.96 38.33 56.38 35.66 42.41 62.67 StarCoder li2023starcoder 28.08 33.95 58.75 22.54 31.83 50.74 25.54 31.45 56.67 StarCoder2-15b lozhkov2024starcoder 27.77 32.60 60.57 18.70 23.28 39.28 23.27 29.78 53.49 Mixtral-8x7B-v0.1 jiang2024mixtral 22.82 29.14 55.90 19.38 25.25 47.22 20.54 26.30 53.40 Phi-2 javaheripi2023phi 19.04 24.56 48.22 14.54 20.34 40.85 14.82 20.69 44.54 Phi-1 gunasekar2023textbooks 14.99 18.38 43.17 12.48 15.42 37.45 12.54 15.96 38.75 Inst-models WizardCoder-Python-13B-V1.0 luo2023wizardcoder 34.31 40.06 62.90 30.99 36.75 59.50 32.54 38.34 64.67 Phind-CodeLlama-34B-v2 30.08 33.49 59.47 25.25 29.40 50.73 27.55 31.85 58.93 CodeLlama-13b-Instruct roziere2023code 28.56 32.67 57.09 26.25 30.72 49.48 26.73 33.50 54.53 GPT-3.5 27.27 37.69 63.59 23.15 33.94 52.79 22.59 33.63 55.22 DeepSeek-Coder-7b-Instruct guo2024deepseek 25.18 29.91 58.50 20.23 26.02 45.76 22.20 27.74 56.69 Mixtral-8x7B-Instruct-v0.1 jiang2024mixtral 23.41 28.71 59.83 19.04 24.83 52.75 20.45 25.84 58.17 InstructPrompt Inst-models WizardCoder-Python-13B-V1.0 26.20 30.68 67.32 24.56 30.25 67.54 24.70 29.56 68.34 CodeLlama-13b-Instruct 25.66 30.82 62.04 27.44 34.11 63.33 26.73 32.43 64.85 GPT-3.5 23.82 39.10 40.55 19.62 36.03 37.48 19.00 34.00 35.28 Mixtral-8x7B-Instruct-v0.1 18.11 23.04 67.73 18.54 23.12 69.66 15.38 20.61 68.86
@@
    Pretrained LLMs produce code that functions correctly but do not effectively reuse the provided dependencies, occasionally duplicating the implementation from the given context. This leads to redundancy and potentially creates technical debt or code smell issues (maldonado2015detecting; sierra2019survey; rasool2015review; santos2018systematic).

### v2

# RepoExec: Evaluate Code Generation with a Repository-Level Executable Benchmark

@@ Abstract
The ability of CodeLLMs to generate executable and functionally correct code at the repository-level scale remains largely unexplored. We introduce RepoExec, a novel benchmark for evaluating code generation at the repository-level scale. RepoExec focuses on three main aspects: executability, functional correctness through automated test case generation with high coverage rate, and carefully crafted cross-file contexts to accurately generate code. Our work explores a controlled scenario where developers specify necessary code dependencies, challenging the model to integrate these accurately. Experiments show that while pretrained LLMs outperform instruction-tuned models in correctness, the latter excel in utilizing provided dependencies and demonstrating debugging capabilities. We also introduce a new instruction-tuned dataset that focuses on code dependencies and demonstrate that CodeLLMs fine-tuned on our dataset have a better capability to leverage these dependencies effectively. RepoExec aims to provide a comprehensive evaluation of code functionality and alignment with developer intent, paving the way for more reliable and applicable CodeLLMs in real-world scenarios. The dataset and source code can be found at https://github.com/FSoft-AI4Code/RepoExec.
Figure 1: Data Processing Pipeline of RepoExec
@@ 1 Introduction
Code Large Language Models (CodeLLMs) have emerged as a prominent solution to assist with coding tasks recently Wang et al. (2021, 2023); Feng et al. (2020); Allal et al. (2023); Li et al. (2023); Lozhkov et al. (2024); Guo et al. (2024); Pinnaparaju et al. (2024); Zheng et al. (2024); Roziere et al. (2023); Nijkamp et al. (2022); Luo et al. (2023); Xu et al. (2022); Bui et al. (2023). While CodeLLMs are adept at capturing the immediate context of the code being written, they often do not fully exploit the broader context available within the entire code repository. This limitation stems from the token length restrictions of general LLMs, which also apply to CodeLLMs. Consequently, suggestions generated by these models may not be optimally aligned with the project’s architecture or intended functionality, as they tend to overlook the rich information embedded in related files, class hierarchies, dependencies, and more. Therefore, evaluating the ability of CodeLLMs on code generation at the repository-level scale is crucial to accurately measure their performance in real-world scenarios. There are existing repository-level code generation benchmarks , such as RepoBench Liu et al. (2023b), RepoCoder Zhang et al. (2023b), CrossCodeEval Ding et al. (2023), CoCoMIC Ding et al. (2024), and DevEval Li et al. (2024), Although these benchmarks provide a comprehensive perspective on evaluating code generation at the repository-level scale, they lack 3 major features that prevent them from accurately reflecting real-world coding scenarios. First, code generation in the real world requires the code to be executable, and none of the existing benchmarks focus on this aspect. Second, these benchmarks rely on simple methods to retrieve contexts, such as including all imports and function calls (RepoBench) or simply retrieving similar files as contexts (RepoCoder), which might introduce noise into the contexts and impact the performance of code generation systems on these benchmarks. Third, generating real world code at repository-level scale requires the CodeLLMs to be able to capture contexts (both cross-file and in-file effectively), which may pose several questions such as if the dependencies provided by human are really the dependencies that the LLM need, or how well the CodeLLMs process the dependencies given that lots of depdencies feed into the prompt, do they use all of them or skip some of them?
@@ 2 Related works
Coding-related tasks have been crucial for assessing the performance of Large Language Models (LLMs), with code generation emerging as a primary focus (Chen et al., 2021; Li et al., 2023; Jiang et al., 2024; Touvron et al., 2023; Roziere et al., 2023; Xu et al., 2022; Allal et al., 2023; Nijkamp et al., 2022). Early benchmarks have been introduced to address this issue Yin et al. (2018); Iyer et al. (2018); Manh et al. (2023); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); however, they often had limited scope or employed weak evaluation approaches. Some benchmarks Yin et al. (2018); Iyer et al. (2018); Manh et al. (2023) exhibit domain diversity akin to real-world applications; however, their evaluation methodologies are constrained to match-based metrics, thereby decreasing the reliability of these benchmarks Chen et al. (2021). Meanwhile, benchmarks with reliable evaluation approaches such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021) and APPS Hendrycks et al. (2021) often entail limitations to specific domains like competitive programming. Besides, all the benchmarks mentioned primarily focus on standalone function generation, lacking consideration for cross-contextual and dependency invocation scenarios.
Several recent studies have introduced frameworks and benchmarks for repository-level code generation (Ding et al., 2024; Shrivastava et al., 2023; Ding et al., 2023; Liao et al., 2023; Liu et al., 2023b). These studies closely align with real-world scenarios, underscoring the importance of cross-contextual considerations. However, these works are still limited to match-based evaluation methods. A recent study Li et al. (2024) introduced the DevEval benchmark for code generation within repository contexts, evaluating performance based on extracted tests available in the repository. These benchmarks primarily focus on assessing the functional correctness of generated outputs but have not extensively investigated the correctness in conjunction with the ability to utilize dependency contexts.
@@ 3.1 Data Source
Developing an executable benchmark within repository contexts is challenging due to complex setup requirements and frequent lack of clear installation guidelines in repositories. Previous studies (Ding et al., 2024; Shrivastava et al., 2023; Ding et al., 2023; Liao et al., 2023) have often relied on match-based metrics for evaluation, which may not fully capture code functionality. In addition, test cases are essential for assessing code functionality. However, extracting test cases from repositories (Zhang et al., 2023a; Li et al., 2024; Zhang et al., 2024) often relies on available functions and heuristic rules, limiting adaptability and excluding data without existing tests. For instance, Li et al. 2024 found that over 99% of functions were discarded due to the absence of suitable tests. We propose a dataset collection pipeline to ensure that repositories can build executable environments and that test cases are automatically generated.
@@ 3.2 Functions and Dependencies Extraction
We extract functions and their dependencies from repositories, considering only those suitable for function-level code generation, akin to benchmarks like HumanEval Chen et al. (2021) and MBPP Austin et al. (2021). Using tree-sitter, we parse files into Abstract Syntax Trees (AST) to extract functions, focusing on those with comprehensive docstrings and excluding functions used as entry points or that do not produce verifiable outputs.
@@ 3.3 Test case generation
To overcome the limitation of requiring available tests in the repository for evaluation (as mentioned in Section 3.1), we leverage large language models (LLMs) to generate test cases automatically. Our proposed approach ensures the correctness of created test cases while also controlling and enhancing test coverage. To execute and validate the generated test cases, it is necessary to configure each repository to create an executable environment. Follow Lemieux et al. 2023, we use pipreqs [^1] https://github.com/bndr/pipreqs to identify each project’s dependencies.
@@ 3.3.1 Correctness Control
In this stage, we present the procedure for generating initial test cases and ensure the tests’ correctness. Specifically, we use CodeLlama-13b Roziere et al. (2023) and provide the model with the prompt detailed in Appendix C. After generation, we extract the first 20 assertion statements to construct the test cases. To ensure their correctness, we employ syntax and execution-based filtering methods to exclude low-quality test cases.
@@ 3.3.2 Coverage Enhancement
Weak unit tests may allow incorrect implementations to pass Liu et al. (2023a). To address this, we propose a strategy for enhancing test coverage using LLMs. Given the complexity of this task, requiring a strong understanding of code, we utilize GPT-3.5 to improve the quality of test cases. We provide GPT-3.5 with three prompts (see Appendix C) to handle challenging scenarios, including edge and corner cases. The initial generated tests serve as few-shot examples. We then extract the newly generated tests and ensure their correctness using our methodology from Section 3.3.1. Table 1 shows a coverage improvement of about 4% after enhancement, reaching 96.25%. The performance gap has also increased to over 5% (Appendix B), indicating greater robustness in the evaluation.
@@ 4.1 Data Formatting
We follow Muennighoff et al. 2023 to define 2 types of prompt formats in our evaluation of RepoExec across LLMs: (1) BasePrompt, which concatenates all contexts with the target functions (Figure 2), and (2) InstructPrompt, which includes specific instructions for the LLMs to follow, utilizing two variations as input formats (further details and examples in Appendix D).
@@ 4.2 Dataset Stastistic
Dataset #Samples Repo context Test case Coverage (%) CoNaLA Yin et al. (2018) 500 - CONCODE Iyer et al. (2018) 2,000 - HumanEval Chen et al. (2021) 164 ✓, H 99.43 MBPP Austin et al. (2021) 974 ✓, H 98.48 RepoCoder Zhang et al. (2023a) 373 ✓, P - CrossCodeEval Ding et al. (2023) 2,665 - CoCoMIC Ding et al. (2024) 6,888 - DevEval Li et al. (2024) 1,874 ✓, P - RepoExec 355 ✓, A 92.46    + coverage-enhancement 96.25
@@
Table 1 compares the details of RepoExec with those of existing code generation benchmarks. Benchmarks that exclude execution-based evaluation Yin et al. (2018); Iyer et al. (2018); Ding et al. (2023, 2024) may gather substantial amounts of data; however, they are inadequate for assessing the quality of the generated code. For HumanEval and MBPP, the majority of problems involve standalone functions, which do not reflect real-world scenarios. Besides, benchmarks that rely on human-annotated and pre-existing test cases Chen et al. (2021); Austin et al. (2021); Zhang et al. (2023b); Li et al. (2024) are challenging to scale and control the test coverage. Finally, repository-context benchmarks (RepoCoder Zhang et al. (2023b), CrossCodeEval Ding et al. (2023), CoCoMIC Ding et al. (2024), and DevEval Li et al. (2024)) primarily focus on evaluating retrieval modules. In contrast, our approach emphasizes the generation module’s ability to understand and utilize human-provided dependency contexts.
@@
#No problem #No testcase Avg tokens
Cross-file Total Avg Coverage (%) Avg Prompt Solution
Full context 22.8% 355 96.25 98.38 362.92 78.46
Medium context - - - - 253.05 -
Short context - - - - 179.66 -
@@ 5.1 Evaluation metrics
We use the well-known Pass@k Chen et al. (2021) to measure the functional correctness of generation outputs. Futhermore, to assess the models’ ability to utilize provided dependencies in accordance with human intent, we introduce the Dependency Invocation Rate (DIR). This metric represents the percentage of invoked dependencies out of the total number of dependencies provided. Let $D_{g}$ denote the set of identifiers in the generated output, and $D_{s}$ denote the set of provided dependencies extracted from the solution. The Dependency Invocation Rate (DIR) is calculated as follows:
@@ 5.2 Evaluation Results
Model Full context Medium context Small context pass@1 pass@5 DIR pass@1 pass@5 DIR pass@1 pass@5 DIR BasePrompt Pre-models CodeLlama-34b-Python Roziere et al. (2023) 42.93 49.54 68.85 35.92 42.95 58.15 39.80 45.79 64.23 CodeLlama-13b-Python Roziere et al. (2023) 38.65 43.24 62.26 32.96 38.33 56.38 35.66 42.41 62.67 StarCoder Li et al. (2023) 28.08 33.95 58.75 22.54 31.83 50.74 25.54 31.45 56.67 StarCoder2-15b Lozhkov et al. (2024) 27.77 32.60 60.57 18.70 23.28 39.28 23.27 29.78 53.49 Mixtral-8x7B-v0.1 Jiang et al. (2024) 22.82 29.14 55.90 19.38 25.25 47.22 20.54 26.30 53.40 Phi-2 Javaheripi et al. (2023) 19.04 24.56 48.22 14.54 20.34 40.85 14.82 20.69 44.54 Phi-1 Gunasekar et al. (2023) 14.99 18.38 43.17 12.48 15.42 37.45 12.54 15.96 38.75 Inst-models WizardCoder-Python-13B-V1.0 Luo et al. (2023) 34.31 40.06 62.90 30.99 36.75 59.50 32.54 38.34 64.67 Phind-CodeLlama-34B-v2 30.08 33.49 59.47 25.25 29.40 50.73 27.55 31.85 58.93 CodeLlama-13b-Instruct Roziere et al. (2023) 28.56 32.67 57.09 26.25 30.72 49.48 26.73 33.50 54.53 GPT-3.5 27.27 37.69 63.59 23.15 33.94 52.79 22.59 33.63 55.22 DeepSeek-Coder-7b-Instruct Guo et al. (2024) 25.18 29.91 58.50 20.23 26.02 45.76 22.20 27.74 56.69 Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024) 23.41 28.71 59.83 19.04 24.83 52.75 20.45 25.84 58.17 InstructPrompt Inst-models WizardCoder-Python-13B-V1.0 26.20 30.68 67.32 24.56 30.25 67.54 24.70 29.56 68.34 CodeLlama-13b-Instruct 25.66 30.82 62.04 27.44 34.11 63.33 26.73 32.43 64.85 GPT-3.5 23.82 39.10 40.55 19.62 36.03 37.48 19.00 34.00 35.28 Mixtral-8x7B-Instruct-v0.1 18.11 23.04 67.73 18.54 23.12 69.66 15.38 20.61 68.86
@@
    Pretrained LLMs produce code that functions correctly but do not effectively reuse the provided dependencies, occasionally duplicating the implementation from the given context. This leads to redundancy and potentially creates technical debt or code smell issues (Maldonado and Shihab, 2015; Sierra et al., 2019; Rasool and Arshad, 2015; Santos et al., 2018).